The manager of a coffee shop has one type of coffee that sells for $6 per pound and another type that sells for $13 per pound. The manager wishes to mix 30 pounds of the $13 coffee to get a mixture that will sell for $9 per pound. How many pounds of the $6 coffee should be used?
Answer by oberobic(2304)

x = lb of one coffee
13x = value of that coffee
.
30-x = lb of other coffee
6(30-x) = value of that coffee
.
30 = lb of coffee in final mix
9(30) = 270 = value of the final mix
.
13x + 6(30-x) = 270
13x + 180 -6x = 270
7x = 90
x = 90/7 lb
x = 12 6/7 lb of the more expensive coffee
.
That means you will add
30-x = 17 1/7 lb of the cheaper coffee
.